integrate the following integral
∫ sin^2(2x +5) dx

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

I = \int sin^{2}(2x+5)dx
I = \frac{1}{2}\int 2sin^{2}(2x+5)dx
I = \frac{1}{2}\int (1-cos(4x+10))dx
I = \frac{1}{2}\int dx-\frac{1}{2}\int cos(4x+10)dx
I = \frac{x}{2}-\frac{1}{2}\int cos(4x+10)dx
I = \frac{x}{2}-\frac{1}{8}sin(4x+10)+c